Please help me solve the problem 3-(-1). Please respond as soon as possible. Thank you.
Sophia W.

Found 2 solutions by plourde216, ivy12003:
Answer by plourde216(7)

The easiest way to complete a problem using a minus sign and a negative is keep-switch-change.
3-(-1) keep the 3 the way it is, switch the minus sign to a plus, change the negative to a positive
3+(+1)=4
You can use the keep-switch-change method in a number of problems but this is where I see it used most often

Answer by ivy12003(22)

When we have two negative signs together, it also means add. (Or, when subtracting integers, you change the subraction sign to a positive and the sign on the right to its reverse sign.) Therefore, the question would be:
3-(-1)= 3 + 1 = 4